# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 1. Introduction
Let $\mathbb{K}$ be a finite field of characteristic $p$ and order $q$, and $G=\mathrm{GL}_{2}(q)$. For $n\geq 1$, define $V_{n}(q)$ to be the set of homogeneous polynomials over $\mathbb{K}$ of degree $n$ in two commuting variables. Then $V_{n}(q)$ becomes a $\mathbb{K}G$-module of dimension $n+1$ with respect to the natural polynomial action of $\mathrm{GL}_{2}(q)$. The modules $V_{n}(q)$ are irreducible precisely when $10\leqslant n\leqslant p-$ 1. Writing $\Lambda(q)$ for the dual module of $V_{p}(q)$, define:
@@
Note that $|S_{n}(q)|=q|V_{n}(q)|=q^{n+2}$ and $|S_{\Lambda}(q)|=q|\Lambda(q)|=q^{p+2}$. In this paper, we refer to the $p$-groups $S_{n}(q)$ and $S_{\Lambda}(q)$ as polynomial $p$-groups. The primary objective of this work is to determine up to isomorphism all core-free saturated fusion systems on $S_{\Lambda}(q)$ and $S_{n}(q)$ when $11\leqslant n\leqslant p-$ 1.
In [CP10], Clelland and the second author introduced two infinite families of saturated fusion systems on the groups $S_{n}(q)$ with $11\leqslant n\leqslant p-$ 1 which can be viewed as generalisations of the $p$-fusion systems of $\operatorname{SL}_{3}(q)$ and, when $q$ is odd, $\mathrm{Sp}_{4}(q)$. Indeed, these groups have parabolic subgroups which are closely related to $P_{1}(q)$ and $P_{2}(q)$ respectively, and the Clelland–Parker fusion systems possess an essential subgroup normaliser of the form $P_{n}(q)$. Other than in some very small cases, the fusion systems exhibited in [CP10] are exotic (see [CP10, Theorems 5.1 and 5.2]). In Section 4, we describe these systems and variants given by adding “field automorphisms” and by removing essential subgroups via a process known as pruning (see [PS21, Lemma 6.4]). Some of these examples had previously been considered in [HLL23]. Collectively, we refer to the fusion systems so obtained as polynomial fusion systems.
Suppose $S=S_{n}(q)$. When $n=1$, $S$ is a Sylow $p$-subgroup of $\operatorname{SL}_{3}(q)$ and core-free saturated fusion systems on $S$ were classified by Clelland in his PhD thesis [Cle07], generalising the pioneering work of Ruiz and Viruel [RV04] when $q=p$ (see Proposition 7.1). When $q$ is odd and $n=2$, $S$ is a Sylow $p$-subgroup of $\mathrm{Sp}_{4}(q)$, and core-free fusion systems on $S_{2}(q)$ were studied by Henke–Shpectorov in the unpublished manuscript [HS]. Apart from the polynomial fusion systems, they discover two exotic fusion systems when $q=9$ related to a certain $3$-dimensional $\mathbb{F}_{9}$-module for $2^{.}\mathrm{PSL}_{3}(4)$ which we refer to as Henke-Shpectorov fusion systems. Our first main theorem is a classification of core-free fusion systems on $S_{n}(q)$ for all $11\leqslant n\leqslant p-$ 1.
@@
Suppose that $p$ is a prime, $q$ is a power of $p$ and $11\leqslant n\leqslant p-$ 1. If $\mathcal{F}$ is a saturated fusion system on $S=S_{n}(q)$ with $O_{p}(\mathcal{F})=1$ then either $\mathcal{F}$ is a polynomial fusion system or one of the following holds:
@@
    $q=p$ and $S$ has an abelian subgroup of index $p$.
@@
    $\mathcal{F}$ is a Henke-Shpectorov fusion system on $S_{2}(9)$.
In particular, all saturated fusion systems on $S_{n}(q)$ are known.
@@
This gives a characterisation of the Henke–Shpectorov systems as the only exceptional core-free exotic fusion systems on polynomial $p$-groups which occur when $q>p$. Theorem 1.1 subsumes the results in [Cle07] and [HS] mentioned above, but we rely on the former since the structure of $\mathrm{Aut}(S_{n}(q))$ when $n=1$ differs from the case $n\geq 2$, and this precludes from a uniform analysis of this group. We do not rely on the results in [HS] but it is notable that, in contrast to [HS], we make more liberal use of the Classification of Finite Simple Groups in our analysis to gain insight into the groups which possess a strongly $p$-embedded subgroup (in [HS] this result is only used to prove that certain fusion systems are exotic). For the conclusion that $\mathcal{F}$ is known in Theorem 1.1(1), we rely on the results of Oliver, and Craven–Oliver-Semeraro [Oli14, COS17]. To sum up, for the proof of Theorem 1.1 we may assume that $q>p$ is odd and a broad brush overview of the strategy we employ is as follows. We first determine the groups $X$ with Sylow $p$-subgroup $S_{n}(q)$ which have the property that $O_{p}(X)$ is isomorphic to $V_{n}(q)$ as an $S_{n}(q)$-module. This turns out to be a rather delicate calculation which takes up a good part of Section 5. After that, in Section 6, we determine up to $S_{n}(q)$-conjugacy the candidates for $\mathcal{F}$-essential subgroups in a saturated fusion system on $S$, before assembling those which form compatible systems in Section 7.
@@
In a later paper the authors, along with other collaborators, will demonstrate that the $p$-groups $S_{n}(q)$ do not support core-free saturated fusion systems whenever $n>p$. In addition, in this article we show in Proposition 6.5 that, provided $q>p$, all saturated fusion systems on $S_{p}(q)$ are constrained. Nonetheless, it turns out that the $p$-groups $S_{\Lambda}(q)$ do support a generic family of core-free saturated fusion systems, which we also regard as polynomial. For $q>p$ these examples are all exotic, new to this article and described in Section 3. They arise as $p^{\prime}$-index subsystems of a certain fixed fusion system and are parametrised by $p^{\prime}$-subgroups of $\mathrm{Aut}(\mathbb{K})$; the reader is referred to Section 4 for a fuller discussion of this correspondence. As in the case of the groups $S_{n}(q)$, in a later paper we will show that the analogous $p$-groups constructed from the duals of $V_{n}(q)$ never support core-free saturated fusion systems whenever $n>p$.
@@
    $p$ is odd and either: $q>p$ and $\mathcal{F}$ is a polynomial fusion system with $\Lambda(q)\ntrianglelefteq\mathcal{F}$; or $q=p$ and $S$ has an abelian subgroup of index $p$.
In particular, all saturated fusion systems on $S_{\Lambda}(q)$ are known.
If $p=2$ then $S$ has nilpotency class two and Theorem 1.2(1) follows from [vB24b, Proposition 4.2]. There it is shown, in addition, that $\mathcal{F}$ is realised by $\mathrm{PSp}_{4}(2)\cong\mathrm{Sym}(6)$ when $q=p$. As above, if $q=p$ is odd then $\mathcal{F}$ is known by [Oli14, COS17]. Thus we may assume that $q>p$ is odd and the strategy for the proof of Theorem 1.2 is similar to that described above for Theorem 1.1.
@@
Context: Theorems 1.1 and 1.2 feed into several avenues of research, and our primary application is to a more encompassing project in which the authors, alongside other collaborators, intend to determine all reduced saturated fusion systems $\mathcal{F}$ on $S$ which have an $\mathcal{F}$-essential subgroup which is both abelian and normal in $S$, thereby extending the work in [Oli14, COS17, OR20]. Computational experiments (see, for example, [PS21]) have revealed that the proportion of $p$-groups which support reduced fusion systems is vanishingly small. For this reason, the discovery and classification of families of (block) exotic fusion systems remains an important pursuit, to which the results of this paper contribute (see [AKO11, Questions III.7.4, IV.7.1]). It has already been observed that the Sylow $p$-subgroup of a simple group $G$ is often a fruitful source of exotic examples. In this direction, numerous authors have considered the case when $G$ is a Lie-type group of characteristic $p$ and rank $2$ [RV04, Cle07, PS18, BFM19, MM22, vB23, HS] culminating in [vB24a]. Notably, this endeavour has led to the discovery of several families of exotic fusion systems (see, for example, [RV04, PS18, HS]).
@@
Our notation for groups is mostly standard and follows [Asc86]. For $A,B\leq G$, we define the $n$-fold commutator $[A,B;n]$ by first setting $[A,B;1]:=[A,B]$ and then inductively defining $[A,B;n]=[[A,B;n-1],B]$ for $n>1$. Similarly we mostly adopt standard conventions for fusion systems as found, for example, in [AKO11]. One exception is our use of $\mathcal{E}(\mathcal{F})$ for the set of $\mathcal{F}$-essential subgroups of a saturated fusion system $\mathcal{F}$. Throughout, we will often investigate actions of a $p$-group on one of its abelian sections. For this reason, we will tend to use multiplicative notation for such actions when viewing these sections as part of the $p$-group, and additive notation when viewing them as vector spaces.
Acknowledgements: We are grateful to David Craven for sharing his knowledge of indecomposable $\operatorname{SL}_{2}(q)$-modules, and in particular for his help with Lemmas 5.6 and 5.8. We also thank Ellen Henke, Justin Lynd and Bob Oliver for helpful discussions.
This work resulted from the workshop “Patterns in Exotic Fusion Systems” funded by the Heilbronn Institute for Mathematical Research. The first author is a member of the GNSAGA INdAM research group and kindly acknowledges its support. The third author gratefully acknowledges funding from the UK Research Council EPSRC for the project EP/W028794/1. The fourth author is supported by the Heilbronn Institute for Mathematical Research.
2. The structure of polynomial $p$-groups and their automorphisms
@@
Finally, regarding $A$ as a $\mathbb{F}_{p}$-space, both $D$ and $\mathrm{Aut}(\mathbb{K})$ act on $A$ and we can check that, for $\phi\in\mathrm{Aut}(\mathbb{K})$ and $(\lambda,\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right))\in D$ we have $(\lambda,\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right))^{\phi}=(\lambda^{\phi},\left(\begin{smallmatrix}a^{\phi}&b^{\phi}\\ c^{\phi}&d^{\phi}\end{smallmatrix}\right))\in D$. In particular, we can form the semidirect product $\mathrm{Aut}(\mathbb{K})\ltimes D$ and this group acts $\mathbb{F}_{p}$-linearly on $A$ by acting on the coefficients of the polynomials. Set
$D^{*}=O^{p}(\mathrm{Aut}(\mathbb{K}))\ltimes D=O_{p^{\prime}}(\mathrm{Aut}(\mathbb{K}))\ltimes D.$
and
@@
Consider $A$ as an $\mathbb{F}_{p}D^{\dagger}$-module and observe that the action of $D^{\dagger}$ on $A$ maps polynomials of degree $k$ to polynomials of degree $k$. For $n\geq 0$, let $V_{n}(q)$ be the $\mathbb{F}_{p}D^{\dagger}$-submodule of $A$ consisting of the homogeneous polynomials of degree $n$. We have $\dim_{\mathbb{K}}V_{n}(q)=n+1$ and provided that $11\leqslant n\leqslant p-$ 1 these submodules are irreducible self-dual $\mathbb{F}_{p}D^{\dagger}$-modules (see [BN41, p588]).
@@
$B=\{\left(\phi,\lambda,\left(\begin{smallmatrix}a&0\\ c&d\end{smallmatrix}\right)\right)\mid a,d,\lambda\in\mathbb{K}^{*},c\in\mathbb{K},\phi\in O^{p}(\mathrm{Aut}(\mathbb{K}))\},$
and
@@
Then $U\in\operatorname{Syl}_{p}(D^{*})$, $U$ is elementary abelian of order $q$, $B=N_{D^{*}}(U)$ has order $q(q-1)^{3}m_{p^{\prime}}$ where $m_{p^{\prime}}=|O^{p}(\mathrm{Aut}(\mathbb{K}))|$ is the $p^{\prime}$-part of $m$ for $q=p^{m}$, and $B=U\Sigma$. Fix $n\in\mathbb{N}$ and define
$P_{n}(q)=D\ltimes V_{n}(q)$
and
$P_{n}^{*}(q)=D^{*}\ltimes V_{n}(q)$
with the action of $D$ and $D^{*}$ as described above. Let
$S_{n}(q)=U\ltimes V_{n}(q)$
be identified as a subgroup of $P_{n}^{*}(q)$. Thus $S_{n}(q)$ is a Sylow $p$-subgroup of $P_{n}(q)$ and $P_{n}^{*}(q)$.
From now we set $P=P_{n}(q)$, $P^{*}=P_{n}^{*}(q)$, $S=S_{n}(q)$, $V=V_{n}(q)$ and we identify $D$, $B$, $\Sigma$, $U$ and $V$ as subgroups of $P^{*}$. This means that we use multiplicative notation when considering $V$ as a $p$-subgroup of $S$, but additive notation when calculating inside $V$ as a vector space. Hence we have $S=UV$ and $O^{p^{\prime}}(P^{*}/V)=O^{p^{\prime}}(P/V)=O^{p^{\prime}}(P)/V\cong\operatorname{SL}_{2}(\mathbb{K})$.
@@
$i{j-1\choose k}=(k+i){j\choose k}\pmod{p}.$
@@
$(x^{i}y^{j}\cdot\left(\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}\right))\psi$ $=$ $(x^{i}(x+y)^{j})\psi=\left(\sum_{k=0}^{j}{j\choose k}x^{k+i}y^{j-k}\right)\psi$
$=$ $\sum_{k=0}^{j-1}{j\choose k}(k+i)x^{k+i-1}y^{j-k-1}=x^{i-1}\sum_{k=0}^{j-1}(k+i){j\choose k}x^{k}y^{j-k-1}$
$=$ $ix^{i-1}\sum_{k=0}^{j-1}{j-1\choose k}x^{k}y^{j-1-k}=ix^{i-1}(x+y)^{j-1}=(x^{i}y^{j})\psi\cdot\left(\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}\right).$
@@
Set ¯=$\overline{\mathcal{B}}=\{\overline{v}\mid v\in\mathcal{B}\}$. Then $\overline{\mathcal{B}}$ is a basis for $\Lambda(q)$. As usual, the action of $D^{*}$ on $\Lambda(q)$ is given by
@@
From now on write $\Lambda=\Lambda(q)$ and $P_{\Lambda}^{*}(q)=D^{*}\ltimes\Lambda$. We identify $\Lambda$ as a subgroup of $P_{\Lambda}^{*}(q)$ and set $S_{\Lambda}(q):=U\Lambda$, a Sylow $p$-subgroup of $P_{\Lambda}^{*}(q)$.
@@
and has order $q-1$. In addition, $C_{D^{*}}(\Lambda(q))=C_{D^{*}}(V_{p}(q))$.
@@
If $G:=\operatorname{SL}_{2}(q)$ with $G\leq\mathrm{Aut}(V_{n}(q))$ and $T\in\mathrm{Syl}_{p}(G)$ then $|C_{G}(C_{V_{n}(q)}(T))|=q.(n,q-1)$.

### v2

@@ 1. Introduction
For a prime $p$, the analysis of $p$-local structure in finite groups and their modular representation theory is naturally expressed in the language of saturated fusion systems on $p$-groups. Since $p$-groups of a given order are overwhelmingly dominated by those of nilpotency class $2$, the work of the fourth author [VAN24b] implies that almost all $p$-groups admit no reduced saturated fusion system. In fact, the proportion of $p$-groups supporting a reduced saturated fusion system appears to be vanishingly small; this expectation is reinforced by computational evidence (see, for example, [PS21]). Consequently, recognizing the $p$-groups that do support reduced fusion systems, and classifying the fusion systems that occur on them, is an interesting and difficult problem. The results of this paper contribute to this programme (see [AKO11, Questions III.7.4, IV.7.1]).
A saturated fusion system on a finite pp–group $S$ is said to be exotic if it does not arise as the fusion system of any finite group with Sylow $p$-subgroup $S$. As we shall see, the fusion systems studied in this paper are predominantly exotic.
One source of $p$-groups which support an exotic fusion system are the Sylow $p$-subgroups of non-abelian simple groups. In particular, a substantial body of work has examined the saturated fusion systems on a $p$-group $S$ which is the Sylow $p$-subgroup of a Lie type group of characteristic $p$ and rank $2$ [RV04, CLE07, PS18, BFM19, MOR22, VAN23, 18], culminating in the classification obtained in [VAN24a]. This direction of research has produced several families of exotic fusion systems (see, for example, [RV04, PS18, 18]). For $q$ a power of $p$, the $p$-groups considered here, which we christen polynomial $p$-groups, are generalisations of the Sylow $p$-subgroups of $\operatorname{SL}_{3}(q)$ and $\mathrm{Sp}_{4}(q)$ and have their origins in polynomial representations of $\operatorname{SL}_{2}(q)$. Thus they lie directly in the rank $2$ Lie type landscape. Moreover, polynomial $p$-groups are known to support reduced fusion systems and therefore provide a natural source of $p$-groups for the general programme described above. From a different perspective, the polynomial $p$-groups denoted $S_{n}(q)$ below are remarkably like “$q$-versions” of maximal class groups (see Lemma 2.8). Thus this work also complements the work in [GP25].
In this article, we provide a complete classification of saturated fusion systems $\mathcal{F}$ on polynomial $p$-groups with $O_{p}(\mathcal{F})=1$ (core-free) and thus a determination of all the reduced saturated fusion systems on such groups. Our key results, Theorems 1.1 and 1.3, form an important step toward a broader project carried out with collaborators in which we determine all reduced saturated fusion systems $\mathcal{F}$ on a $p$-group $S$ that possess an $\mathcal{F}$-essential subgroup that is both abelian and normal in $S$. This project extends the work initiated in [OLI14, COS17, OR20], and the classification obtained here supplies input required for the principal families of fusion systems arising in that extension.
Let $\mathbb{K}$ be a finite field of characteristic $p$ and order $q$, and $G=\mathrm{GL}_{2}(q)$. For $n\geq 1$, define $V_{n}(q)$ to be the set of homogeneous polynomials over $\mathbb{K}$ of degree $n$ in two commuting variables. Then $V_{n}(q)$ becomes a $\mathbb{K}G$-module of dimension $n+1$ with respect to the natural polynomial action of $\mathrm{GL}_{2}(q)$. The modules $V_{n}(q)$ are irreducible precisely when $0\leqslant n\leqslant p-1$. Writing $\Lambda(q)$ for the dual module of $V_{p}(q)$, define:
@@
Note that
$|S_{n}(q)|=q|V_{n}(q)|=q^{n+2}$
and
$|S_{\Lambda}(q)|=q|\Lambda(q)|=q^{p+2}.$
In this paper, we refer to the $p$-groups $S_{n}(q)$ and $S_{\Lambda}(q)$ as polynomial $p$-groups. The primary objective of this work is to determine up to isomorphism all core-free saturated fusion systems $\mathcal{F}$ on $S_{\Lambda}(q)$ and $S_{n}(q)$ when $1\leqslant n\leqslant p-1$.
In [CP10], Clelland and the second author introduced two infinite families of saturated fusion systems on the groups $S_{n}(q)$ with $1\leqslant n\leqslant p-1$ which can be viewed as generalisations of the $p$-fusion systems of $\operatorname{SL}_{3}(q)$ and, when $q$ is odd, $\mathrm{Sp}_{4}(q)$. Indeed, these groups have parabolic subgroups which are closely related to $P_{1}(q)$ and $P_{2}(q)$ respectively, and the Clelland–Parker fusion systems possess an essential subgroup normaliser of the form $P_{n}(q)$. Other than in some very small cases, the fusion systems exhibited in [CP10] are exotic (see [CP10, Theorems 5.1 and 5.2]). In Section 4, we describe these systems and variants given by adding “field automorphisms” and by removing essential subgroups via a process known as pruning (see [PS21, Lemma 6.4]). Some of these examples had previously been considered in [HLL23]. Collectively, we refer to all of these fusion systems as polynomial fusion systems.
Suppose $S=S_{n}(q)$. When $n=1$, $S$ is a Sylow $p$-subgroup of $\operatorname{SL}_{3}(q)$ and core-free saturated fusion systems on $S$ were classified by Clelland in his PhD thesis [CLE07], generalising the pioneering work of Ruiz and Viruel [RV04] when $q=p$ (see Proposition 7.1). When $q$ is odd and $n=2$, $S$ is a Sylow $p$-subgroup of $\mathrm{Sp}_{4}(q)$, and core-free fusion systems on $S_{2}(q)$ were studied by Henke–Shpectorov in the unpublished manuscript [18]. Apart from the polynomial fusion systems, they discover two exotic fusion systems when $q=9$ related to a certain $6$-dimensional $\mathbb{F}_{3}$-module for $2^{.}\mathrm{PSL}_{3}(4)$ which we refer to as Henke–Shpectorov fusion systems. Our first main theorem is a classification of core-free fusion systems on $S_{n}(q)$ for all $1\leqslant n\leqslant p-1$.
@@
Suppose that $p$ is a prime, $q$ is a power of $p$ and $1\leqslant n\leqslant p-1$. If $\mathcal{F}$ is a saturated fusion system on $S=S_{n}(q)$ with $O_{p}(\mathcal{F})=1$ then either $\mathcal{F}$ is a polynomial fusion system or one of the following holds:
@@
    $q=p$ and $S$ has an elementary abelian subgroup of index $p$.
@@
    $\mathcal{F}$ is a Henke–Shpectorov fusion system on $S_{2}(9)$.
In particular, all reduced saturated fusion systems on $S_{n}(q)$ are known.
@@
This gives a characterisation of the Henke–Shpectorov systems as the only exceptional core-free exotic fusion systems on polynomial $p$-groups which occur when $q>p$. Theorem 1.1 subsumes the results in [CLE07] and [18] mentioned above, but we rely on the former since the structure of $\mathrm{Aut}(S_{n}(q))$ when $n=1$ differs from the case $n\geq 2$, and this precludes from a uniform analysis of this group. We do not rely on the results in [18] but it is notable that, in contrast to [18], we make more liberal use of the Classification of Finite Simple Groups in our analysis to gain insight into the groups which possess a strongly $p$-embedded subgroup (in [18] this result is only used to prove that certain fusion systems are exotic). For the conclusion that $\mathcal{F}$ is known in Theorem 1.1(1), we rely on the results of Oliver, and Craven–Oliver-Semeraro [OLI14, COS17]. To sum up, for the proof of Theorem 1.1 we may assume that $q>p$ is odd and a broad brush overview of the strategy we employ is as follows. We first determine the groups $X$ with Sylow $p$-subgroup $S_{n}(q)$ which have the property that $O_{p}(X)$ is isomorphic to $V_{n}(q)$ as an $S_{n}(q)$-module. This turns out to be a rather delicate calculation which takes up a good part of Section 5. After that, in Section 6, we determine up to $S_{n}(q)$-conjugacy the candidates for $\mathcal{F}$-essential subgroups in a saturated fusion system on $S$, before assembling those which form compatible systems in Section 7.
@@
In a later paper the authors, along with other collaborators, will demonstrate that the $p$-groups $S_{n}(q)$ do not support core-free saturated fusion systems whenever $n>p$. In addition, in this article we show in Proposition 6.5 that, provided $q>p$, all saturated fusion systems on $S_{p}(q)$ are constrained. Nonetheless, it turns out that the $p$-groups $S_{\Lambda}(q)$ do support a generic family of core-free saturated fusion systems, which we also regard as polynomial.
Theorem 1.2.
Suppose that $p$ is an odd prime, and $q=p^{n}>p$. Then $S_{\Lambda}(q)$ supports a simple, exotic fusion system $\mathcal{F}_{\Lambda}(q)$.
The system $\mathcal{F}_{\Lambda}(q)$ is described in Section 3, where Theorem 1.2 is a consequence of Theorem 3.9. In Remark 3.10, it is proved that these examples do not coincide with currently known exotic families whenever $q>p$. We include the case $q=p$ in our construction in Section 3, but note that these examples were already considered in [COS17]. The systems described in Theorem 1.2 arise as $p^{\prime}$-index subsystems of a certain fixed fusion system $\mathcal{F}_{\Lambda}^{*}(q)$, also described in Section 3. Indeed, these subsystems are parameterised by $p^{\prime}$-subgroups of $\mathrm{Aut}(\mathbb{K})$, and so we obtain several new exotic systems this way. The reader is referred to Section 4 for a fuller discussion of this correspondence. For $n>p$, as for the groups $S_{n}(q)$, we will show in a later paper that the analogous $p$-group associated to the dual module of $V_{n}(q)$ also does not support core-free saturated fusion systems. We thus focus our attention on the group $S_{\Lambda}(q)$ when $n=p$. Our third main result classifies all core-free fusion systems on this group.
@@
    $p$ is odd and either:
    1. (a)
      $q>p$ and $\mathcal{F}$ is a polynomial fusion system with $\Lambda(q)$ not normal in $\mathcal{F}$; or
    2. (b)
      $q=p$ and $S$ has an elementary abelian subgroup of index $p$.
In particular, all reduced saturated fusion systems on $S_{\Lambda}(q)$ are known.
If $p=2$ then $S$ has nilpotency class two and Theorem 1.3(1) follows from [VAN24b, Proposition 4.2]. There it is shown, in addition, that $\mathcal{F}$ is realised by $\mathrm{PSp}_{4}(2)\cong\mathrm{Sym}(6)$ when $q=p$. As above, if $q=p$ is odd then $\mathcal{F}$ is known by [OLI14, COS17]. Thus we may assume that $q>p$ is odd and the strategy for the proof of Theorem 1.3 is similar to that described above for Theorem 1.1.
@@
Our notation for groups is mostly standard and follows [ASC86]. For $A,B\leq G$, we define the $n$-fold commutator $[A,B;n]$ by first setting $[A,B;1]:=[A,B]$ and then inductively defining [A,B;n]=[[A,B;n1],B][A,B;n]=[[A,B;n-1],B] for $n>1$. Similarly we mostly adopt standard conventions for fusion systems as found, for example, in [AKO11]. One exception is our use of $\mathcal{E}(\mathcal{F})$ for the set of $\mathcal{F}$-essential subgroups of a saturated fusion system $\mathcal{F}$. Throughout, we will often investigate actions of a $p$-group on one of its abelian sections. For this reason, we will tend to use multiplicative notation for such actions when viewing these sections as part of the $p$-group, and additive notation when viewing them as vector spaces.
Acknowledgements: We are grateful to David Craven for sharing his knowledge of indecomposable $\operatorname{SL}_{2}(q)$-modules, and in particular for his help with Lemmas 5.6 and 5.8. We also thank Ellen Henke, Justin Lynd and Bob Oliver for helpful discussions. We wish to the thank the anonymous referee whose comments greatly improved the exposition of this work.
This work resulted from the workshop “Patterns in Exotic Fusion Systems” funded by the Heilbronn Institute for Mathematical Research. The first author is a member of the GNSAGA INdAM research group and thankfully acknowledges its support. The third author gratefully acknowledges funding from the UK Research Council EPSRC for the project EP/W028794/1. The fourth author is supported by the Heilbronn Institute for Mathematical Research.
2. The structure of polynomial $p$-groups and their automorphism group
@@
Finally, regarding $A$ as a $\mathbb{F}_{p}$-space, both $D$ and $\mathrm{Aut}(\mathbb{K})$ act on $A$ and we can check that, for $\phi\in\mathrm{Aut}(\mathbb{K})$ and $(\lambda,\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right))\in D$ we have $(\lambda,\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right))^{\phi}=(\lambda^{\phi},\left(\begin{smallmatrix}a^{\phi}&b^{\phi}\\ c^{\phi}&d^{\phi}\end{smallmatrix}\right))\in D$. In particular, we can form the semidirect product $\mathrm{Aut}(\mathbb{K})\ltimes D$ and this group acts $\mathbb{F}_{p}$-linearly on $A$ by acting on the coefficients of the polynomials.
Notation 2.1.
Set
$D:=\mathbb{K}^{*}\times\mathrm{GL}_{2}(\mathbb{K}),$
$D^{*}:=O^{p}(\mathrm{Aut}(\mathbb{K}))\ltimes D=O_{p^{\prime}}(\mathrm{Aut}(\mathbb{K}))\ltimes D,\mbox{ and }$
@@
Consider $A$ as an $\mathbb{F}_{p}D^{\dagger}$-module and observe that the action of $D^{\dagger}$ on $A$ maps polynomials of degree $k$ to polynomials of degree $k$. For $n\geq 0$, let $V_{n}(q)$ be the $\mathbb{F}_{p}D^{\dagger}$-submodule of $A$ consisting of the homogeneous polynomials of degree $n$. We have $\dim_{\mathbb{K}}V_{n}(q)=n+1$ and provided that $1\leqslant n\leqslant p-1$ these submodules are irreducible self-dual $\mathbb{F}_{p}D^{\dagger}$-modules (see [BN41, p588]). We introduce notation to describe particular subgroups of $D^{*}$.
Notation 2.2.
@@
$B=\{\left(\phi,\lambda,\left(\begin{smallmatrix}a&0\\ c&d\end{smallmatrix}\right)\right)\mid a,d,\lambda\in\mathbb{K}^{*},c\in\mathbb{K},\phi\in O^{p}(\mathrm{Aut}(\mathbb{K}))\},\mbox{ and }$
@@
Then $\in\operatorname{Syl}_{p}(D^{*}), UU$ is elementary abelian of order $q$, and $B=N_{D^{*}}(U)$ has order $q(q-1)^{3}m_{p^{\prime}}$ where $m_{p^{\prime}}=|O^{p}(\mathrm{Aut}(\mathbb{K}))|$ is the $p^{\prime}$-part of $m$ for $q=p^{m}$, and $B=U\Sigma$. The following groups will be of primary importance:
Notation 2.3.
Fix $n\in\mathbb{N}$. With the action of $D$ and $D^{*}$ as described above, define
$P_{n}(q)=D\ltimes V_{n}(q),\hskip 5.69054ptP_{n}^{*}(q)=D^{*}\ltimes V_{n}(q),\hskip 5.69054pt\mbox{and}\hskip 5.69054ptS_{n}(q)=U\ltimes V_{n}(q).$
Thus $S_{n}(q)$ is a Sylow $p$-subgroup of $P_{n}(q)$ and $P_{n}^{*}(q)$. We identify $D$, $B$, $\Sigma$, $U$ and $V$ as subgroups of $P_{n}(q)^{*}$. This means that we use multiplicative notation when considering $V$ as a $p$-subgroup of $S$, but additive notation when calculating inside $V$ as a vector space. Hence we have $S_{n}(q)=UV_{n}(q)$ and $O^{p^{\prime}}(P_{n}^{*}(q)/V_{n}(q))=O^{p^{\prime}}(P_{n}(q)/V_{n}(q))=O^{p^{\prime}}(P_{n}(q))/V_{n}(q)\cong\operatorname{SL}_{2}(\mathbb{K})$.
@@
$i\binom{j-1}{k}=(k+i)\binom{j}{k}\pmod{p}$.
@@
$(x^{i}y^{j}\cdot\left(\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}\right))\psi$ $=$ $(x^{i}(x+y)^{j})\psi=\left(\sum_{k=0}^{j}\binom{j}{k}x^{k+i}y^{j-k}\right)\psi$
$=$ $\sum_{k=0}^{j-1}\binom{j}{k}(k+i)x^{k+i-1}y^{j-k-1}=x^{i-1}\sum_{k=0}^{j-1}(k+i)\binom{j}{k}x^{k}y^{j-k-1}$
$=$ $ix^{i-1}\sum_{k=0}^{j-1}\binom{j-1}{k}x^{k}y^{j-1-k}=ix^{i-1}(x+y)^{j-1}=(x^{i}y^{j})\psi\cdot\left(\begin{smallmatrix}1&0\\ 1&1\end{smallmatrix}\right).$
@@
Set $\overline{\mathcal{B}}=\{\overline{v}\mid v\in\mathcal{B}\}$. Then $\overline{\mathcal{B}}$ is a basis for $\Lambda(q)$. As usual, the action of $D^{*}$ on $\Lambda(q)$ is given by
@@
Notation 2.5.
Define:
$P_{\Lambda}^{*}(q)=D^{*}\ltimes\Lambda\mbox{ and }S_{\Lambda}(q):=U\ltimes\Lambda(q).$
Thus $S_{\Lambda}(q)$ is a Sylow $p$-subgroup of $P_{\Lambda}^{*}(q)$. Moreover, from now on, we write $\Lambda=\Lambda(q)$ and identify $\Lambda$ as a subgroup of $P_{\Lambda}^{*}(q)$.
@@
and has order q1q-1. In addition, $C_{D^{*}}(\Lambda(q))=C_{D^{*}}(V_{p}(q))$.
@@
If $X:=\operatorname{SL}_{2}(q)$ with $X\leq\mathrm{Aut}(V_{n}(q))$ and $T\in\mathrm{Syl}_{p}(X)$ then $|C_{X}(C_{V_{n}(q)}(T))|=q.(n,q-1)$.